OpenStudy (anonymous):

How do I find (gof)(-3) when f=2x+1 and g=x^2-1? (o is a composition symbol)

OpenStudy (anonymous):

gof(x)=(2x+1)^2-1 = 4x^2+4x now, gof(-3)=4*(-3)^2+4*(-3)=24

OpenStudy (anonymous):

\[f(-3)=f(-3)+1=-6+1=-5\] \[g(-5)=(-5)^2-1=25-1=24\]

OpenStudy (anonymous):

should have written \[f(-3)=2\times -3+1=-5\] sorry

OpenStudy (anonymous):

I like Satellite73 way better? Why the difference?

OpenStudy (anonymous):

\[g\circ f(x)\] means first do f, then do g

OpenStudy (anonymous):

they are the same really

OpenStudy (anonymous):

another typos \[g\circ f(x)=g(f(x))=g(2x+1)=(2x+1)^2-1\]

OpenStudy (anonymous):

basically, you substitute (with parentheses) what f is equal to namely (2x+1) into g = x^2-1... so that looks like (2x+1)^2-1 , and put -3 into that.. so {2(-3)+1}^2-1